a shopkeeper claims that the average life of CFL bulb is 1600 hours.to check this claim a researcher takes a sample of 100 CFL bulbs of tha same make randomly and finds mean lifetime of 1570 hours with standard daviation of 120 hours.is tha claim acceptable at 5% level of significance?


1
Expert's answer
2021-09-21T09:12:02-0400

"n=100 \\

\bar{x}=1570 \\

\sigma= 120 \\

H_0: \mu=1600 \\

H_1: \mu≠1600"

Test-statistic:

"Z = \frac{\bar{x} - \mu }{\sigma/ \sqrt{n}} \\

= \frac{1570-1600}{120/ \sqrt{100}} \\

= -2.5 \\

α=0.05 \\"

Two-tailed test.

Reject H0 if Z ≤1.96 or Z≥1.96.

"Z=-2.5 < Z_c=-1.96"

We are able to reject H0.

The population mean lifetime of bulbs is not equal to 1600 hours at 0.05 significance level.